The perimeter of a square is 3 yards what is the length of one side of the square in inches?
Answer by jim_thompson5910(35256)

P = 4*s Perimeter of square formula (P is perimeter, s is side length)


3 = 4*s Plug in the given info


3/4 = s


s = 3/4


So each side is 3/4 or 0.75 inches long